# The Neglected Tails of Vision-Language Models

### 23 Jan 2024 v1

@@ Abstract
Vision-language models (VLMs) excel in zero-shot recognition but exhibit drastically imbalanced performance across visual concepts. For example, CLIP, despite an impressive mean zero-shot accuracy on ImageNet (72.7%), yields $<$10% on ten concepts (e.g., gyromitra and night snake), presumably because these concepts are under-represented in VLMs’ imbalanced pretraining data. Yet, assessing this imbalance is challenging as it is non-trivial to calculate the frequency of specific concepts within VLMs’ large-scale pretraining data. Our work makes the first attempt to measure the concept frequency by analyzing pretraining texts. We use off-the-shelf language models to help count relevant texts that contain synonyms of the given concepts and resolve linguistic ambiguity. We confirm that popular VLM datasets like LAION indeed exhibit long-tailed concept distributions, which strongly correlate with per-class accuracies. Further, contemporary multimodal systems, e.g., visual chatbots and text-to-image generators, also struggle with the rare concepts identified by our method. To mitigate VLMs’ imbalanced performance in zero-shot recognition, we propose REtrieval-Augmented Learning (REAL). First, instead of prompting VLMs using the original class names, REAL uses their most frequent synonyms found in VLMs’ pretraining texts. This already outperforms human-engineered and LLM-generated prompts over nine benchmark datasets, likely because VLMs have seen more images associated with the frequently used synonyms. Second, REAL uses all the concept synonyms to retrieve a small, class-balanced set of pretraining data to train a robust classifier. REAL surpasses the recent retrieval-augmented solution REACT, using 400$\times$ less storage and 10,000$\times$ less training time!
@@ 1 Introduction
Figure 1: Vision-language models (VLMs) inherit long tails from their pretraining data. (a) VLMs’ pretraining datasets, such as LAION-400M [35] and LAION-2B [36], exhibit long-tailed distributions for visual concepts defined in downstream tasks like ImageNet [9]. We sort the 1K ImageNet classes according to their frequency in LAION-400M calculated with our concept frequency estimation method (cf. Fig. 2). (b) For zero-shot recognition, OpenCLIP models [14] trained on LAION-400M and LAION-2B respectively yield per-class accuracies that strongly correlate with the long-tailed concept frequencies (binned on a log-scale). Interestingly, other VLMs such as CLIP [33] and MetaCLIP [47] (trained on private data) also show similar imbalanced performances, likely because their web-scraped pretraining datasets follow similar long-tailed distributions of the real world. (c) Our method helps identify rare concepts, such as the night snake, which is one of the most tailed ImageNet concepts. We show that state-of-the-art multimodal systems, including GPT4-V [48], LLaVA [21], DALL-E 3 [39], and SD-XL [2], all fail to recognize or generate it. The supplement contains more examples.
@@
Vision-language models (VLMs) such as CLIP [33] play a pivotal role in mainstream multimodal applications, including visual chatbots [21, 29] and text-to-image generation [34, 3]. Their efficacy largely derives from their large-scale image-text pretraining datasets like LAION [35, 36] that encompass broad visual concepts.
@@
Imbalanced performance of VLMs. Despite their impressive visual understanding capabilities, VLMs exhibit an imbalanced performance in downstream tasks. For example, in the established zero-shot recognition task where no training examples are available, CLIP [33] yields a high mean accuracy (72.7%) across 1K semantic classes on ImageNet [9] but only $<$10% accuracy on ten classes (e.g., gyromitra and night snake) (Fig. 1b). This motivates us to delve into the imbalanced performance of VLMs (e.g., CLIP [33] and OpenCLIP [14]) through the lens of zero-shot image recognition.
Why do VLMs exhibit imbalanced performance? A plausible reason is their inherent bias [50], which originates from the imbalanced distribution of visual concepts in their pretraining data. However, to date, the literature has not established a clear link between VLMs’ imbalanced performance and the concept distribution in their training datasets, presumably because estimating concept frequency in such large-scale, multimodal pretraining data is non-trivial.
@@
Concept frequency estimation. Estimating class frequency in a typical classification dataset is straightforward, where counting class occurrences in annotated labels is sufficient. But estimating concept frequency in VLMs’ pretraining datasets is non-trivial, as they contain free-form texts with significant lexical variation, e.g., a sneaker can also be referred to as a running shoe or a trainer. To address this, we leverage off-the-shelf large language models (LLMs). Fig. 2 illustrates our approach, which begins by asking an LLM (such as ChatGPT [29]) to enumerate all synonyms for a specific concept. We then use string matching to find all pretraining texts that contain the concept or its synonyms. However, due to linguistic ambiguity, the initially retrieved texts can include unrelated phrases, such as “tiger shark in water” for the target concept tiger (a mammal). We again use an LLM to filter out irrelevant texts to the target concept and only count relevant ones for its frequency measure. Our simple method for calculating concept frequency allows us to, for the first time, unveil three key insights: (1) it exposes long-tailed concept distributions in VLMs’ pretraining data (Fig. 1a); (2) it establishes a correlation between the long-tailed distribution and VLMs’ imbalanced performance (Fig. 1b); and (3) it sheds light on why the cutting-edge foundational models (e.g., GPT4-Vision in Fig. 1c) fail to recognize images featuring rare concepts. Importantly, our analysis provides insights that facilitate the exploration of approaches (explained below) to enhance VLMs in downstream tasks such as zero-shot image recognition.
@@
Improving VLMs’ zero-shot performance. Motivated by our analysis, we present REtrieval-Augmented Learning (REAL) to address the biased predictions of VLMs, thereby greatly improving VLMs’ zero-shot recognition performance. REAL has two variants. First, observing that some synonyms appear more frequently in VLM’s pretraining data than the original concept names, we propose REAL-Prompt. It replaces the original concept names with their most frequent synonyms for zero-shot recognition. For example, we substitute cash machine with ATM, as the latter appears ten times more frequent in LAION (Fig. 3). This simple change already surpasses previous methods that use either human-engineered prompts [33] or LLM-generated ones [26, 32] (cf. Table 1). We attribute its success to VLMs’ greater exposure to these more frequent synonyms, which improve their accuracy in recognizing related images. Second, inspired by retrieval-augmented strategies [17, 10, 22, 18, 42], we introduce REAL-Linear, which reuses relevant pretraining data to better adapt VLMs for zero-shot recognition without using data from downstream tasks. The key idea is to retrieve a small, class-balanced set of images from pretraining data to train a robust linear classifier [20]. In contrast to prior arts [22, 42] that perform costly feature-based retrieval by running VLMs to compute image or text features, our method implements text-based retrieval via string matching, achieving a significant boost in efficiency. As a result, our REAL resoundingly outperforms the recent retrieval-augmented solution, REACT [22], using 400$\times$ less storage and 10,000$\times$ less training time (cf. Table 1 and 3)!
@@
  • We propose a method for estimating the frequency of visual concepts within VLMs’ large-scale pretraining data. Our analysis, for the first time, exposes long-tailed concept distributions in popular datasets like LAION and reveals systematic failures of VLMs [48, 39, 21, 2] in handling rare concepts.
  • To address the biased predictions of VLMs, we propose REAL to improve their zero-shot performance without using data from downstream tasks. REAL establishes a new state-of-the-art in zero-shot recognition through its efficient prompting and retrieval-augmented training strategies.
@@ 2 Related Works
Biases in foundation VLMs. Pretrained on expansive multimodal datasets [4, 43, 7], VLMs often exhibit biases related to gender, race, and geography [25], leading to imbalanced predictions in downstream applications [1, 27]. Such biases likely stem from data imbalances in their pretraining set [50]. Recent studies [45, 50, 37] seek to mitigate imbalanced predictions of VLMs by training on additional data from downstream tasks. Despite these efforts, there has been a lack of analysis of the imbalances in the pretraining data itself. Our study presents the first direct analysis of VLMs’ pretraining data, uncovering a long-tailed distribution of concepts and its strong correlation with the imbalanced performance of VLMs. Our analysis tool can also help researchers identify rare, inadequately learned concepts in VLMs, reducing the risk of biased predictions.
@@
Prompting VLMs for zero-shot recognition. VLMs demonstrate remarkable capabilities in zero-shot recognition tasks, where only the names of target concepts are provided without corresponding training images. CLIP [33] shows that ensembling concept names with human-engineered prompt templates, such as “ a photo of a {class}” and “a demonstration of a {class}”, often enhances zero-shot recognition. Approaches like DCLIP [26] and CuPL [32] employ LLMs (e.g., GPT-3) to create class-specific prompts with visual descriptions, for example, “a tiger, which has sharp claws”. While most works focus on refining prompt templates [23, 38, 49], they use the provided class names as is. A recent work [30] suggests that prompting with common English names instead of Latin scientific names improves zero-shot recognition of fine-grained species. Differently, our method proposes to replace the original class names with their most common synonyms found in the pretraining dataset. This simple change already outperforms existing methods that strive to improve prompt templates. Moreover, our approach can be integrated with existing methods to further improve their zero-shot recognition performance.
@@
Retrieval-augmented strategy. The natural language processing (NLP) literature introduces this strategy to address challenging tasks such as knowledge-based question-answering [10, 17]. For example, in order to produce a better answer given a question, this strategy retrieves relevant facts from an external knowledge base (e.g., the Internet or pretraining dataset) to ground LLMs on the most accurate, up-to-date information. To improve zero-shot visual recognition, recent works [22, 18, 42] propose to finetune VLMs on relevant images retrieved from VLMs’ pretraining datasets for better accuracy. While solutions such as REACT [22] are effective across diverse target domains, they require significant computing resources, such as hundreds of GPU hours and intensive memory usage for large-batch contrastive training. In contrast to these works, we propose a lightweight solution that retrieves data by fast string matching of concept synonyms in pretraining texts, bypassing the heavy computation of VLM features. Moreover, we learn a robust, parameter-efficient linear classifier atop the frozen VLM, enhancing training efficiency. Our solution not only outperforms existing computationally intensive approaches, but also offers an opportunity for research on retrieval-augmented strategy with modest compute.
@@
Figure 2: Using large language models (LLMs) to estimate concept frequency in a VLM’s pretraining dataset. We conduct the frequency estimation using publicly available LAION [35] datasets. First, since a visual concept can be expressed in various ways, we ask an LLM (e.g., ChatGPT [29]) to enumerate all its synonyms to search for all relevant pretraining texts. For example, for tiger, we retrieve all captions that contain not only “tiger” but also its synonyms such as “Panthera tigris”, using efficient string matching. Second, we filter out irrelevant captions that do not refer to the target concept by its definition. For example, although “tiger shark swimming in water” contains “tiger”, it actually refers to a type of shark, not the animal tiger as defined by “Panthera tigris, a large, striped Asian cat”. We conduct the filtering process by querying an LLM (cf. Section 3).
@@ 3 The Long-Tailed Concept Distribution
This section describes our approach for estimating concept frequency in VLM’s pretraining data, followed by a discussion of our key findings.
@@ 3.1 Concept Frequency Estimation
The foremost challenge in estimating concept frequency is the sheer size of a VLM’s pretraining dataset. For example, the popular open-source LAION-400M [36] dataset (used for training OpenCLIP [14]) takes $\sim$10TB of physical space. We instead estimate concept frequency directly from pretraining texts, bypassing the need for pretraining images. This enables us to download just the text metadata, requiring only  60GB space for LAION-400M. Our method consists of the following two steps (cf. Fig. 2).
Step 1: Deriving synonyms for concepts of interest. A well-known issue in NLP is lexical variation, meaning that a concept can be expressed in multiple ways. For example, “sneaker” can be referred to as “running shoes” or “athletic footwear”, and “tiger” may also be called “Panthera tigris”. To account for the lexical variation in concept frequency estimation, we first derive a list of synonyms [^1] For simplicity, we use the term “synonyms” in a broader sense to encompass all forms of lexical variation, including but not limited to traditional synonyms, idiomatic expressions, and different phrasings that convey the same or similar meanings. for a given visual concept. To do so, we turn to an off-the-shelf LLM (e.g., ChatGPT [29]), by asking a simple question “What are some common ways of referring to {concept}?”. With all the synonyms, we retrieve relevant pretraining texts using string matching. The string matching process is quite efficient — it takes only 5 hours to retrieve 400M pretraining texts from LAION-2B containing ImageNet’s 1K concepts.
Figure 3: Demonstration of REAL-Prompt, which replaces the given concept names (e.g., “cash machine”) with their most frequent synonyms (e.g., “ATM”) in the prompt template, e.g., “a photo of {concept}”. REAL-Prompt uses an LLM (ChatGPT) to obtain synonyms for a given concept, followed by string matching to identify the most frequently occurring ones in pretraining texts. We demonstrate REAL-Prompt on some ImageNet concepts with their most frequent synonyms, frequencies (in LAION-400M), and per-class accuracies. The simple change in prompts significantly improves zero-shot recognition. We detail the procedure for REAL-Prompt in Sec. 4.1 and compare against prior works in Table 1.
Step 2: Filtering out irrelevant pretraining texts. Using simple string matching to retrieve pretraining texts may be inaccurate due to linguistic ambiguities. For instance, the concept “tiger”, defined as “Panthera tigris, a large, striped Asian cat”, might appear in unrelated contexts in the retrieved captions such as “tiger shark swimming in water” and “Tiger Woods, a famous golf player”; in these texts, the “tiger” actually refers to a shark species and a celebrity, respectively. To tackle these ambiguities at an affordable cost, we utilize the state-of-the-art open-source LLM Llama-2 [40]. For each retrieved text, we ask:
@@
In real-world applications, concepts are typically defined in the labeling policies of downstream tasks. In this work, to align with standard benchmarks, we adopt definitions from Elevater [19]. Finally, captions identified as irrelevant to the target concept by the LLM are excluded from retrieved data.

### 02 Feb 2024 v2

@@ Abstract
Vision-language models (VLMs) excel in zero-shot recognition but their performance varies greatly across different visual concepts. For example, although CLIP achieves impressive accuracy on ImageNet (60-80%), its performance drops below 10% for more than ten concepts like night snake, presumably due to their limited presence in the pretraining data. However, measuring the frequency of concepts in VLMs’ large-scale datasets is challenging. We address this by using large language models (LLMs) to count the number of pretraining texts that contain synonyms of these concepts. Our analysis confirms that popular datasets, such as LAION, exhibit a long-tailed concept distribution, yielding biased performance in VLMs. We also find that downstream applications of VLMs, including visual chatbots (e.g., GPT-4V) and text-to-image models (e.g., Stable Diffusion), often fail to recognize or generate images of rare concepts identified by our method. To mitigate the imbalanced performance of zero-shot VLMs, we propose REtrieval-Augmented Learning (REAL). First, instead of prompting VLMs using the original class names, REAL uses their most frequent synonyms found in pretraining texts. This simple change already outperforms costly human-engineered and LLM-enriched prompts over nine benchmark datasets. Second, REAL trains a linear classifier on a small yet balanced set of pretraining data retrieved using concept synonyms. REAL surpasses the previous zero-shot SOTA, using 400$\times$ less storage and 10,000$\times$ less training time!
@@ 1 Introduction
Figure 1: Vision-language models (VLMs) inherit long tails from their pretraining data. (a) VLMs’ pretraining datasets, such as LAION-400M [37] and LAION-2B [38], exhibit long-tailed distributions for visual concepts defined in downstream tasks like ImageNet [10]. We sort the 1K ImageNet classes according to their frequency in LAION-400M calculated with our concept frequency estimation method (cf. Fig. 2). (b) For zero-shot recognition, OpenCLIP models [16] trained on LAION-400M and LAION-2B respectively yield per-class accuracies that strongly correlate with the long-tailed concept frequencies (binned on a log-scale). Interestingly, other VLMs such as CLIP [35] and MetaCLIP [49] (trained on private data) also show similar imbalanced performances, likely because their web-scraped pretraining datasets follow similar long-tailed distributions of the real world. (c) Our method helps identify rare concepts, such as the night snake, which is one of the most tailed ImageNet concepts. We show that state-of-the-art multimodal systems, including GPT-4V [50], LLaVA [23], DALL-E 3 [41], and SD-XL [2], all fail to recognize or generate it. The supplement shows more examples.
@@
Vision-language models (VLMs) such as CLIP [35] play a pivotal role in mainstream multimodal systems, including visual chatbots [23, 31] and text-to-image generation [36, 3]. Their efficacy largely stems from their web-scale image-text pretraining datasets like LAION [37, 38] that cover a wide range of visual concepts.
@@
Imbalanced performance of VLMs. Despite their strong capabilities in visual understanding, VLMs often exhibit biased performance in downstream tasks. For instance, in zero-shot visual recognition tasks (which do not use training samples), CLIP [35] achieves up to 80% mean accuracy across 1K semantic classes on ImageNet [10] but less than $<$10% on specific classes such as night snake (Fig. 1b). This motivates us to explore why VLMs are imbalanced, a crucial yet ever-neglected issue.
Why do VLMs exhibit imbalanced performance? It is commonly believed that foundational VLMs inherit biases [27] from web-scale pretraining data. However, we find no direct evidence linking VLMs’ imbalanced performance to the concept distribution in pretraining data, likely because there is no such tool for measuring concept frequency in large multimodal datasets like LAION [37].
@@
Concept frequency estimation. Estimating class frequency in a typical classification dataset is straightforward, where counting class occurrences using annotated labels is sufficient. However, estimating concept frequency in VLMs’ pretraining datasets is more complex because their free-form texts (or captions) contain significant lexical variations, e.g., a sneaker might also be called a running shoe or a trainer. To address this, we leverage off-the-shelf large language models (LLMs). Fig. 2 illustrates our approach, which begins by asking an LLM (such as ChatGPT [31]) to enumerate all synonyms for a specific concept. We then use string matching to find all pretraining texts that contain the concept or its synonyms. However, due to linguistic ambiguity, the initially retrieved texts may contain irrelevant phrases, such as “tiger shark in water” for the target concept tiger (a mammal). We again use an LLM to filter out such irrelevant texts, and conduct human studies to verify the accuracy of our frequency measure. Our method for calculating concept frequency unveils three key insights: (1) it confirms that VLMs’ pretraining data is indeed long-tailed (Fig. 1a); (2) it shows VLMs perform better on well-represented concepts and worse on under-represented ones (Fig. 1b); and (3) it explains why recent multimodal systems (e.g., GPT-4Vision and DALL-E 3 in Fig. 1c) struggle with rare concepts. Our analysis also provides technical insights to counteract the bias in VLMs, leading to state-of-the-art zero-shot performance in downstream tasks such as image recognition.
@@
State-of-the-art zero-shot recognition. Motivated by our frequency estimation, we introduce REtrieval-Augmented Learning (REAL) to mitigate biased performance of zero-shot VLMs. REAL has two variants. First, observing that some synonyms are more frequent in VLM’s pretraining data than the original concept names, we propose REAL-Prompt. Specifically, we replace given concept names with their most frequent synonyms. For example, cash machine is replaced with ATM, which is ten times more frequent in LAION (Fig. 3). This minor change already surpasses costly human-engineered [35] and LLM-enriched prompts like CuPL [34] (cf. Table 1). Second, inspired by retrieval-augmented strategies [19, 12, 24, 20, 44], we introduce REAL-Linear, which reuses relevant pretraining data to better adapt VLMs without using data from downstream tasks. The key idea is to retrieve a small, balanced set of images from pretraining data to train a robust linear classifier [22]. In contrast to prior arts [24, 44] that perform costly feature-based retrieval by running VLMs to compute image or text features, our method implements text-based retrieval via string matching, achieving a significant boost in efficiency. As a result, our REAL resoundingly outperforms the recent retrieval-augmented SOTA, REACT [24], using 400$\times$ less storage and 10,000$\times$ less training time (cf. Table 1 and 3)!
@@
  • We propose a method for estimating the frequency of visual concepts in VLMs’ large-scale pretraining data. Our analysis, for the first time, exposes long-tailed concept distributions in popular datasets like LAION and reveals systematic failures of VLMs [50, 41, 23, 2] in handling rare concepts.
  • We propose REAL to address the biased performance of zero-shot VLMs. REAL establishes a new state-of-the-art in zero-shot recognition through its efficient prompting and retrieval-augmented training strategies.
@@ 2 Related Works
Biases in foundation VLMs. Pretrained on large-scale multimodal datasets [4, 45, 8], VLMs often exhibit biases related to gender, race, and geography [27], leading to imbalanced predictions in downstream tasks [1, 29]. Recent studies [47, 52, 39] seek to mitigate imbalanced predictions of VLMs by training on additional data from downstream tasks. Despite these efforts, there is no analysis of the imbalances within the pretraining data itself. Our study presents the first examination of VLMs’ pretraining datasets, revealing a long-tailed distribution of concepts that closely correlates with VLMs’ imbalanced performance. Our analytical tool also identifies rare concepts that VLMs have insufficiently learned, thereby preventing biases in downstream applications.
@@
Prompting VLMs for zero-shot recognition. VLMs excel in zero-shot recognition tasks, where only the names of target concepts are provided without corresponding training images. CLIP [35] shows that putting given concept names in human-engineered prompt templates, such as “a photo of a {class}” and “a demonstration of a {class}”, often enhances zero-shot recognition. LLM-enriched approaches like DCLIP [28] and CuPL [34] create class-specific prompts by appending rich visual descriptions generated by LLM, for example, “a tiger, which has sharp claws”. While most works focus on refining prompt templates [25, 40, 51], they use the provided class names as is. A recent work [32] suggests that prompting with common English names instead of Latin scientific names improves zero-shot recognition of fine-grained species. Differently, our REAL-Prompt replaces given class names with their most common synonyms found in the pretraining texts. This simple change outperforms existing methods with much less ChatGPT querying costs. Moreover, our approach can be combined with existing prompt templates to further improve performance.
@@
Retrieval-augmented strategy. Introduced in the natural language processing (NLP) literature, this strategy addresses challenging tasks such as knowledge-based question-answering [12, 19] by retrieving relevant facts from an external knowledge source (e.g., the Internet or a pretraining dataset) to ground LLMs on the most accurate, up-to-date information. To improve zero-shot visual recognition, recent works [24, 20, 44] finetune VLMs on images retrieved from VLMs’ pretraining datasets. While methods like REACT [24] are effective, they demand significant computing resources, including hundreds of GPU hours and extensive memory for large-batch contrastive finetuning. In contrast, our REAL-Linear uses fast string matching to retrieve data based on concept synonyms, thus avoiding costly VLM-based feature extraction. Moreover, we train a parameter-efficient linear classifier atop the frozen VLM [22], which significantly enhances efficiency. Our method not only sets a new state-of-the-art in zero-shot recognition, but also opens avenues for retrieval-augmented research within a modest computational budget.
@@
Figure 2: Using large language models (LLMs) to estimate concept frequency in a VLM’s pretraining dataset. We conduct the frequency estimation using publicly available LAION [37] datasets. First, since a visual concept can be expressed in various ways, we ask an LLM (e.g., ChatGPT [31]) to enumerate all its synonyms to search for potentially relevant pretraining texts. For example, for tiger, we retrieve all captions that contain not only “tiger” but also its synonyms such as “Panthera tigris”. Second, we filter out irrelevant captions that do not refer to the target concept by its definition. For example, although “tiger shark swimming in water” contains “tiger”, it actually refers to a type of shark, not the animal tiger as defined by “Panthera tigris, a large, striped Asian cat”. We conduct the filtering process by querying an LLM Llama-2 [42] (cf. Section 3).
@@ 3 The Long-Tailed Concept Distribution
This section outlines our approach for estimating concept frequency in VLM’s pretraining data and presents our key findings from the analysis.
@@ 3.1 Concept Frequency Estimation
The foremost challenge in estimating concept frequency is the sheer size of a VLM’s pretraining dataset. For example, the popular open-source LAION-400M [38] dataset (used for training OpenCLIP [16]) takes $\sim$10TB of physical storage space. Instead, we estimate concept frequency directly from pretraining texts, eliminating the need to download images. This allows us to download only the text metadata, requiring only $\sim$60GB space for LAION-400M. Next, we proceed with the following two steps (cf. Fig. 2).
Step 1: Deriving synonyms for target concepts. A well-known issue in NLP is lexical variation, meaning that a concept can be expressed in multiple ways. For example, “sneaker” can be referred to as “running shoes” or “athletic footwear”, and “tiger” may also be called “Panthera tigris”. To account for lexical variation, we first derive a list of synonyms [^1] For simplicity, we use the term “synonyms” in a broader sense to encompass all forms of lexical variation, including but not limited to traditional synonyms, idiomatic expressions, and different phrasings that convey the same or similar meanings. for a given visual concept. To do so, we turn to an off-the-shelf LLM (e.g., ChatGPT [31]), by querying a simple question “What are some common ways of referring to {concept}?”. Then, we use string matching to retrieve all pretraining texts containing these synonyms. This matching process is remarkably efficient — it takes only 5 hours to retrieve 400M pretraining texts from the LAION-2B dataset for ImageNet’s 1K concepts. Importantly, these derived concept synonyms can also be used to improve downstream zero-shot recognition, which we discuss in Sec. 4.
Figure 3: Demonstration of REAL-Prompt, which replaces the given concept names (e.g., “cash machine”) with their most frequent synonyms (e.g., “ATM”) in the prompt template, e.g., “a photo of {concept}”. REAL-Prompt uses an LLM (ChatGPT) to obtain a list of synonyms for a given concept, followed by string matching to identify the most frequently occurring ones in pretraining texts. We demonstrate REAL-Prompt on some ImageNet concepts with their most frequent synonyms, frequencies (in LAION-400M), and per-class accuracies (OpenCLIP ViT-B/32). The simple name change in prompts significantly improves zero-shot recognition. We detail the procedure for REAL-Prompt in Sec. 4.1 and compare against prior works in Table 1.
Step 2: Filtering out irrelevant pretraining texts. Using simple string matching may be inaccurate because it could retrieve irrelevant captions due to linguistic ambiguities. For instance, the concept “tiger”, defined as “Panthera tigris, a large, striped Asian cat”, might appear in irrelevant contexts in the retrieved captions such as “tiger shark swimming in water” and “Tiger Woods, a famous golf player”. In these retrieved texts, the “tiger” actually refers to a shark species and a celebrity, respectively. To tackle these ambiguities at an affordable cost, we utilize the state-of-the-art open-source LLM Llama-2 [42]. For each retrieved text, we ask:
@@
In real-world applications, concepts are typically defined in the labeling policies of downstream tasks [6, 11]. In this work, to align with standard benchmarks, we adopt definitions from Elevater [21]. Finally, retrieved captions that are identified as irrelevant to the target concept by the LLM are excluded. We only count the remaining retrieved text to estimate concept frequency.